OpenStudy (gigi1103):

What is the solution for the system of equations? Use the substitution method to solve. 2x+y=20−5 y=−6x+12

OpenStudy (gigi1103):

@3mar

OpenStudy (3mar):

Well, I am here. That is a nice question and easy at the same time!

OpenStudy (gigi1103):

wait its 2x+y=20 -5y= -6x+12

OpenStudy (3mar):

We can use substitution method to plug one equation into the other one!

OpenStudy (3mar):

To do so: - solve for one variable (say y) in one equation. - plug this equation into the other equation. Can you show me how do you understand it?

OpenStudy (gigi1103):

OH

OpenStudy (gigi1103):

i get that but i didnt give me anything to plug in

OpenStudy (3mar):

Can you share your works/steps, please?

OpenStudy (gigi1103):

ok

OpenStudy (gigi1103):

ok so 2*-6+-5 = 20 ? right

OpenStudy (3mar):

It would better to show me how you did the math, @gigi1103

OpenStudy (gigi1103):

ok i was just making sure i set it up right

OpenStudy (3mar):

Ok, Take your time.

OpenStudy (gigi1103):

so 2* - 6 = -12 + -5 = -17

OpenStudy (gigi1103):

so thats not right is it?

OpenStudy (3mar):

You started with which equation?

OpenStudy (gigi1103):

the first one

OpenStudy (gigi1103):

2x+y=20-5

OpenStudy (3mar):

Great! So you solved for y?

OpenStudy (gigi1103):

OH

OpenStudy (gigi1103):

yea ill do the other one now

OpenStudy (3mar):

wait

OpenStudy (3mar):

when you solved the first equation for y, what you got?

OpenStudy (gigi1103):

-17

OpenStudy (3mar):

You did not get me! what I mean is: \[2x+y=20\] \[2x+y-2x=20-2x\] \[y=20-2x\] that is what we need! got it??

OpenStudy (gigi1103):

OH subtracting from both sides

OpenStudy (mathmale):

In other words, 3mar has solved 2x+y=20 for y. This is what you have to do. Substituting (20-2x) for y in the other equation will now elimiinate y; you end up with an equation in x only.

OpenStudy (3mar):

\(\Large y=20-2x\) this is what we will plug it into the other equation!

OpenStudy (gigi1103):

so for the other one

OpenStudy (3mar):

Yes, to have y in one side "alone" and the other terms in the other side!

OpenStudy (3mar):

Just replace \(y\) in the other equation with this equation [\(\Large y=20-2x\)]

OpenStudy (gigi1103):

wait im confused

OpenStudy (gigi1103):

do i plug in the 2x next to the -6

OpenStudy (3mar):

I will simplify it for you!